How high should a 2 kg mass be lifted from the ground if it is thrown upward at speed of 15 m/s? ( Assume g=10 m/s²)


1
Expert's answer
2022-02-14T12:58:26-0500

Given:

"m=2\:\rm kg"

"v_0=15\:\rm m/s"


The maximum height that mass reaches

"h_{\max}=\frac{v_0^2}{2g}=\rm \frac{(15\: m/s)^2}{2*10\: m/s^2}=11.25\: m"